15. Medication order reads: DSNS 4 L over 24 hrs. Drop factor is 15 gtts/mL. Flow rate= gtt(s)/min
Answer #1

15.Given

Order:D5NS

Total Volume:4L(4000ml)

Drop factor:15gtt/ml

Time:24hours

Flow rate:Total volume (ml)×Drop factor ÷Time(hr)×60

=4000×15÷24×60

=4000÷24×4

=4000÷96

=41.6 or 42

Ans:42gtt/min